OpenStudy (anonymous):

3x – 7y = 5 9y = 5x + 5

OpenStudy (anonymous):

3x - 7y = 5 9y = 5x + 5 9y - 5x = 5 both are equal to five , equate 3x - 7y = 9y - 5x

OpenStudy (anonymous):

solve it :D

OpenStudy (anonymous):

A. (–3, 2) B. (8, 5) C. (–10, –5) D. (0, 0)

OpenStudy (anonymous):

any of these work?